#1. A project in its 26th week has an actual cost of $270,000. It was scheduled to have spent $261,000. For the work performed to date, the budgeted value is $272,000. What are the cost and schedule variances for the project? What are the SPI and CPI? What is your assessment of the project from a cost and schedule perspective?

#2. A project has just completed the 87th item in its plan. It was scheduled to have spent $168,000 at this point in the plan, but has actually spent only $156,000. The foreman estimates that the value of the work actually finished is about $162,000. What are the spending and schedule variances for the project? What are the SPI and CPI? What is your assessment of the project from a cost and schedule perspective?

#3. The following project is at the end of its sixth week. Find the cost and schedule variances. Also find the CPI and SPI. Then find the critical ratio of the project using earned value calculations.

Activity

Predecessors

Duration (wks)

Budget ($)

Actual Cost ($)

% Complete

a

-

2

300

400

100

b

-

3

200

180

100

c

a

2

250

300

100

d

a

5

600

400

20

e

b,c

4

400

200

20
Answer #1

1. Actual cost (AC) = 270,000

Earned value (EV) = 272000

Planned value (budget (PV) = 261000

Cost variance = EV (earned value) – AC (actual cost) = $272,000 - $270,000 = $2,000

Schedule variance = EV – PV (planned cost) = $272,000 - $261,000 = $11,000

CPI (cost performance index) = EV/AC =$272,000 / $270,000 = 1.0074

SPI (schedule performance index) = EV/ PV =$272,000 / $261,000 = 1.0421

Both CV and SV are positive i.e. project is running faster than schedule and at a lower cost. CPI & SPI Ratios >1 confirm the same.

2. Actual cost (AC) = 156000

Earned value (EV) =162000

Planned value (budget (PV) = 168000

Cost variance =162000 - 156000 = 6000

Schedule variance = 162000 - 168000 = -6000

CPI = 162000 / 156000 = 1.038

SPI = 162000 / 168000 = 0.964

CV is positive and SV is negative: The project is behind schedule but under budget. Less tasks have been performed than were scheduled at this point, but the tasks that have been performed are at a cost lesser than budgeted.

3. By week 6 , Activity a is completed, b is completed, c is completed, d would have completed 4 weeks and e would have completed 2 weeks.

Planned value = 300 + 200 + 250 + 4/5 * 600 + 1/2 * 400 = 1430 (Note - D has completed 4 weeks and E has two weeks, hence we take proportion of cost)

Earned value = Actual complete * PLanned values = 300 + 200 + 250 + 20%*600 + 20%*400 = 950

Actual cost = 400 + 180 + 300 + 400 +200 =1480

Cost variance = EV (earned value) – AC (actual cost) = 950 - 1480 = - 530

Schedule variance = EV – PV = 950 - 1430 = -480

CPI = 950/1480 = 0.64

SPI = 950 / 1430 = 0.66

Project is behind schedule and over budget.

Critical ratio = SPI * CPI = 0.426